OpenStudy (winsteria):

MEDAL AND FAN Solve. 3x – 6 = −2x + 9 x =?

OpenStudy (danjs):

add 2x to both sides first

OpenStudy (winsteria):

I have my answer, can you check it?

OpenStudy (danjs):

write down all the steps

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

this might help and you can check your work

OpenStudy (danjs):

to do a check of your answer, replace x with the number you got, and see if both sides are the same

OpenStudy (winsteria):

3? @DanJS

OpenStudy (danjs):

3(3) -6 = -2(3)+9 ??

OpenStudy (danjs):

9 - 6 = -6 + 9 3 = 3 true, that is right

OpenStudy (winsteria):

Yay thank you